OpenStudy (anonymous):

5^(-T/2)=.20 I know the answer, but HOW does one arrive there? What rules are in play here?

OpenStudy (anonymous):

@Gralwyn

OpenStudy (turingtest):

helps to notice that 0.2=1/5, so you can rewrite 5^(-T/2)=1/5 now take the log base 5 of both sides

OpenStudy (anonymous):

how does one do that correctly?

OpenStudy (turingtest):

do you know about logarithms?

OpenStudy (anonymous):

\[5^{\frac{ -T }{ 2 }}=0.20=\frac{ 20 }{100 }=\frac{ 1 }{5 }=5^{-1}\] \[\frac{ -T }{2 }=-1,T=?\]

OpenStudy (anonymous):

Yes, but I was under the assumption that I could take the log of just one side so \[(-T/2)\log5{.20}\]

OpenStudy (anonymous):

5 is supposed to be the base there.

OpenStudy (turingtest):

you cannot "just take the log of one side"\[x=y\neq\log x\]

OpenStudy (anonymous):

ok, so log5(-T/2)=log5?

OpenStudy (anonymous):

T=-1*-2=2